the volume of a gas is 400. ml sample of gas at 30.0 c. If the temperature is increased to 50.0 c without changing the pressure, what is the new volume of the gas?
1
Expert's answer
2019-03-29T08:50:31-0400

To solve this problem,we use Gay-Lussac law for isobaric process:


"V1\/T1=V2\/T2"

Convert temperature to kelvin:30 c = 303 K

50 c = 323K

New volume of the gas:


"V2=V1*T2\/T1=400*323\/303=426.4 ml"